A ball of mass m = 0.490 kg is dropped near the surface of the Earth. However, there exists an air resistance F = −βv, where β = 0.700 N · s/m. After a long time, the ball approaches a speed v. What is this speed?


1
Expert's answer
2022-02-18T08:23:52-0500

Given:

"m=0.490\:\rm kg"

"\beta=0.700\:\rm N\cdot s/m"

"g=9.8\:\rm m/s^2"


The terminal velocity of a ball

"v_t=\frac{mg}{\beta}=\frac{0.490*9.8}{0.700}=6.86\:\rm m/s"